# Differentially private federated learning for localized control of infectious disease dynamics

### v1

@@ Discussion
One goal of our study was to explore the possibility of localized predictions at a finer spatial resolution than the county-level. Although we developed a probabilistic approach to disaggregate county data to the community-level, the resulting data proved inadequately representative of actual local dynamics. The random assignment of cases to communities based on population data led to datasets that are unrealistic, particularly for small communities, and potentially do not reflect actual epidemic spread. This experience underscores the challenges of working with sensitive health data at a highly resolved spatial level. While protecting privacy is a key concern, the availability of granular data is crucial for precise localized predictions. Due to the lack of case data resolved on community-level with often only several hundreds of individuals, we could not test our approach on very small population units. Future research will thus consider the creation of a more realistic synthetic dataset. This dataset should follow the structural patterns as observed during the COVID-19 pandemic on community-level. Furthermore, it should be created such that it can be published for scientific research. To do so, we will reach out to public health experts with access to more finely resolved data series.
@@
A limitation of our study concerns the quality and resolution of the underlying data. Clearly, it would have been better if we could also have validated our approach on the much smaller communities. Besides that we attempted to mitigate weekend effects by applying a 7-day moving average, these partially remain and limit the models learning ability. Eventually, problem of underreporting prevails: the reported case numbers represent only a fraction of actual infections with differences in spatial levels. According to data from the Robert Koch Institute [29], test positivity rates in 2022 ranged between 30% and 50%, indicating a substantial number of unreported cases. This underreporting affects the prediction quality, especially if the extent of underreporting varies temporally and spatially.
@@ Acknowledgments
This work was supported by the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics).

### v2

@@ Discussion
One goal of our study was to explore the possibility of localized predictions at a finer spatial resolution than the county-level. Although we developed a probabilistic approach to disaggregate county data to the community-level, the resulting data proved inadequately representative of actual local dynamics. The random assignment of cases to communities based solely on population data likely disrupted local spread patterns, resulting in time series dominated by unstructured noise rather than learnable transmission dynamics. In contrast, real-world community data, while noisy, retains the causal structure of infection chains which FL can leverage. This experience underscores the challenges of working with sensitive health data at a highly resolved spatial level. While protecting privacy is a key concern, the availability of granular data is crucial for precise localized predictions. Due to the lack of case data resolved on community-level with often only several hundreds of individuals, we could not test our approach on very small population units. Future research will thus consider the creation of a more realistic synthetic dataset. This dataset should follow the structural patterns as observed during the COVID-19 pandemic on community-level. Furthermore, it should be created such that it can be published for scientific research. To do so, we will reach out to public health experts with access to more finely resolved data series.
@@
German counties vary substantially in population size (from under 50,000 to over 3.5 million inhabitants), which raises concerns about whether smaller counties might be disadvantaged in a federated learning setting. To address this, we conducted an analysis stratifying model performance by population clusters. The results show that while smaller counties exhibit higher absolute errors due to lower case counts, the privacy-utility tradeoff remains relatively uniform across all population sizes. Importantly, smaller counties benefit from the global model by learning temporal patterns that would be undetectable in their local data alone. The gradient clipping mechanism in our DP-FL algorithm ($S=0.5$) prevents larger counties from dominating the aggregation, ensuring that updates from all counties contribute equally.
We base our experiments on an established client-level DP-FL setup to isolate and interpret the impact of the privacy budget on forecast quality. Prior work has proposed a range of refinements, including local-DP federated learning with per-client privacy modules [28], algorithms that study different locations for adding noise and analyze their convergence and utility guarantees [26], and healthcare-focused FL frameworks [31, 32, 33, 34]. These methods target algorithmic improvements and can in principle be combined with our setting. We expect that these algorithms may further improve forecast accuracy, a systematic evaluation of such variants for infectious disease dynamics should be included in future work. Regarding the model choice, we utilized a standard MLP to demonstrate the impact of the privacy mechanism; however, the framework is compatible with any gradient-based model. We consider this work as a proof of concept with more work required on the way to this methodology being used as designated.
A limitation of our study concerns the quality and resolution of the underlying data. Clearly, it would have been better if we could also have validated our approach on the much smaller communities. Besides that we attempted to mitigate weekend effects by applying a 7-day moving average, these partially remain and limit the models learning ability. Eventually, problem of underreporting prevails: the reported case numbers represent only a fraction of actual infections with differences in spatial levels. According to data from the Robert Koch Institute [39], test positivity rates in 2022 ranged between 30 % and 50 %, suggesting a substantial dark figure. However, we assume that the underreporting rate remains relatively constant within the short timeframes considered. Future research should incorporate auxiliary data sources, such as wastewater-based indicators or dark figure estimators, to better account for true infection numbers.
@@
Supporting information
S1 Tab.
Performance metrics across a fine grid of privacy budgets.
S1 Text.
Formal derivation of the privacy guarantees using the Gaussian Mechanism and Rényi Differential Privacy (RDP). The appendix includes definitions, the privacy accounting methodology, and the conversion theorems used to establish the final $(\varepsilon,\delta)$-differential privacy bounds.
@@ Acknowledgments
This work was supported by the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.